### v1

# Let’s Verify Math Questions Step by Step

Abstract
Large Language Models (LLMs) have recently achieved remarkable progress in mathematical reasoning. To enable such capabilities, many existing works distill strong reasoning models into long chains of thought or design algorithms to construct high-quality math QA data for training. However, these efforts primarily focus on generating correct reasoning paths and answers, while largely overlooking the validity of the questions themselves. In this work, we propose Math Question Verification (MathQ-Verify), a novel five-stage pipeline designed to rigorously filter ill-posed or under-specified math problems. MathQ-Verify first performs format-level validation to remove redundant instructions and ensure that each question is syntactically well-formed. It then formalizes each question, decomposes it into atomic conditions, and verifies them against mathematical definitions. Next, it detects logical contradictions among these conditions, followed by a goal-oriented completeness check to ensure the question provides sufficient information for solving. To evaluate this task, we use existing benchmarks along with an additional dataset we construct, containing 2,147 math questions with diverse error types, each manually double-validated. Experiments show that MathQ-Verify achieves state-of-the-art performance across multiple benchmarks, improving the F1 score by up to 25 percentage points over the direct verification baseline. It further attains approximately 90% precision and 63% recall through a lightweight model voting scheme. MathQ-Verify offers a scalable and accurate solution for curating reliable mathematical datasets, reducing label noise and avoiding unnecessary computation on invalid questions. Our code and data are available at https://github.com/scuuy/MathQ-Verify.
1 Introduction
Large Language Models (LLMs) have demonstrated impressive performance across a wide range of tasks [16, 10, 26]. In particular, when equipped with chain-of-thought (CoT) of varying lengths [36], some models exhibit the ability to solve moderately complex mathematical problems, such as those from the AIME competition [24]. This reasoning capability is largely attributed to high-quality data sources and efficient training frameworks [9].
However, most existing large-scale mathematical question-answering datasets are composed primarily of synthetically generated QA pairs. A critical issue in such datasets is that if the question itself is flawed, such as being ill-posed or logically inconsistent, then the answer cannot be correct—underscoring the fundamental importance of question correctness.
While some recent efforts have focused on verifying the correctness of answers [32, 41, 33], they often neglect to validate the questions themselves, implicitly assuming that all provided problems are well-formed and mathematically sound. This assumption frequently breaks down in real-world or synthetic data scenarios, where generated problems may include internal contradictions or violate foundational mathematical principles [42]. In such cases, the correctness of the question is uncertain, posing a significant challenge for downstream reasoning tasks. They face the following challenges:
(1) Lack of comprehensive question validation methods.
While several recent studies have begun to examine the validity of mathematical questions [19, 17], their focus is typically limited to a narrow set of error types—such as missing assumptions or vague premises. These efforts fall short of establishing a systematic and comprehensive framework for identifying ill-posed or flawed problems. As a result, many QA datasets, particularly those generated synthetically, still include questions that suffer from internal inconsistencies, logical contradictions, or violations of fundamental mathematical principles.
(2) Lack of stepwise and high-difficulty benchmarks for question verification.
Existing benchmarks, such as MathClean [19], do not provide sufficiently challenging questions, nor do they include fine-grained, stepwise annotations needed to evaluate each stage of a multi-step question verification pipeline. This limits the ability to rigorously assess models’ capabilities in detecting and reasoning about complex flaws in mathematical problem formulations.
To systematically address the challenge of question validity, we propose a structured evaluation framework named MathQ-Verify, which comprises five sequential verification stages: (1) Contaminated Instruction Detection, (2) Linguistic Error Detection, (3) Atomic Condition Error Detection, (4) Cross-condition Conflict Detection, and (5) Condition Completeness Validation. These stages collectively ensure a robust and comprehensive assessment of question quality. An ablation study further confirms the necessity and effectiveness of each individual component within the pipeline.
To enable rigorous evaluation of our pipeline, we introduce a stepwise annotated benchmark dataset, ValiMath, which extends MathClean by incorporating more diverse and challenging examples. ValiMath is constructed from filtered synthetic data derived from the NuminaMath dataset [18], consisting of 2,147 math problems sampled from an initial pool of 10,000 automatically generated questions. Each sample is annotated with fine-grained, stepwise validity labels, encompassing both well-formed and flawed cases to support comprehensive model evaluation across all verification stages.
To summarize, our main contributions are summarized as follows:
  • We construct a new dataset ValiMath, by integrating synthetic problems from NuminaMath and enriching them with structured stepwise labels, specifically designed to support comprehensive evaluation of the correctness of mathematical questions.
  • We propose MathQ-Verify, a pipeline that progressively verifies the correctness of math questions by decomposing them into structured components and checking each part against formalized criteria. MathQ-Verify achieves state-of-the-art results on the two evaluation sets of MathClean, and improves F1 by nearly 15% on our ValiMath compared to direct verification baseline.
  • We conduct extensive experiments to assess the effectiveness of each component within the MathQ-Verify pipeline. Through ablation studies, we systematically validate the individual contribution of each verification stage to the overall performance. Additionally, we demonstrate that incorporating a majority voting strategy across verification outputs significantly enhances precision, achieving up to 90%, which highlights the robustness and reliability of our proposed approach.
2 Related Work
LLMs and Mathematical Reasoning.
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of NLP tasks, often surpassing human expert performance in knowledge-intensive domains [7, 21], and generating creative content in writing and design [12, 22]. Beyond general-purpose tasks, recent efforts have increasingly focused on enhancing the reasoning abilities of LLMs, aiming to equip models with deeper, more human-like thinking processes for complex problem solving [28, 39, 31]. Mathematics, as a domain that demands rigorous logical reasoning, has become a popular benchmark for evaluating these capabilities. Commercial products like ChatGPT [25], Claude [4] and Gemini [34] have shown strong performance on math evaluations, while open-source models such as DeepSeek-R1 [9] and the Qwen3 series [5] have also made competitive advances, highlighting the importance of training objectives like reinforcement learning. These advancements have been driven by the availability of high-quality reasoning datasets, especially in math problem solving [20, 8, 2, 14]. Our work enhances this trend by proposing a novel approach to filter and verify math problems, improving data reliability for training and evaluation.
Synthetic Data.
Training LLMs requires massive amounts of data [38, 1], but existing web corpora have become increasingly insufficient to support full training pipelines, particularly in the pretraining stage [35, 29]. In addition, many open-source datasets suffer from noisy supervision and inconsistent quality, making them suboptimal for model training [15]. To address this, recent work has explored synthetic data generation using LLMs [23, 30], enabling controllable data distributions and customizable supervision signals. However, synthetic data especially in the domain of mathematical question generation often contains subtle or critical errors in both questions and solutions, highlighting the need for rigorous verification before such data can be reliably used for training or evaluation. Based on this, we explore methods for verifying and filtering erroneous mathematical data.
Verification in Math QA.
Much attention has been given to improving the accuracy of LLMs on math problem answering, including developing better reasoning strategies and addressing common model failure modes [37, 40]. Models fine-tuned with reinforcement learning have even demonstrated the ability to solve competition-level math problems [9, 25, 6]. However, when presented with flawed or ill-posed questions, LLMs often lack the ability to detect inconsistencies or exhibit overthinking behaviors [27, 11]. Recent efforts such as MathClean [19] and QuestBench [17] have introduced benchmarks for identifying flawed problems and missing assumptions, revealing substantial gaps in model robustness. Despite these advances, structured and comprehensive verification for math problems has yet to receive broad attention. To bridge this gap, our work proposes a progressive, formalized verification pipeline that identifies and filters flawed problems based on a structured decomposition of the problem semantics.
3 Methodology
In this section, we introduce our MathQ-Verify, as shown in Figure 1, designed to verify the correctness of synthetic questions. Our methodology systematically deconstructs each question into its fundamental components and evaluates them step by step through a carefully designed pipeline. This pipeline includes multiple stages, including structural analysis, logical consistency checks, and contextual validation, to ensure a comprehensive assessment of the question’s quality and reasonableness. By rigorously analyzing and verifying each question, our approach aims to establish a robust framework for evaluating the validity and soundness of mathematical problems.
3.1 Problem Formulation
We define the set of questions in the mathematical QA dataset as $\mathcal{Q}=\{q_{i}\}_{i=1}^{L}$, where each $q_{i}$ denotes a natural language math question. This set can be partitioned into two disjoint subsets: a set of well-formed (valid) questions $\mathcal{Q}^{\text{valid}}$ and a set of ill-formed (invalid) questions $\mathcal{Q}^{\text{invalid}}$, such that:
$Q=Q_{\text{valid}}\cup Q_{\text{invalid}},\quad Q_{\text{valid}}\cap Q_{\text{invalid}}=\emptyset$
Identifying whether a given question is well-formed is crucial for downstream tasks such as automated solving and dataset refinement. Therefore, we aim to learn a verification function $V:\mathcal{Q}\rightarrow\{0,1\}$, where $V(q_{i})=1$ indicates that $q_{i}$ is a valid question, and $V(q_{i})=0$ otherwise.
To determine whether a question $q_{i}$ is valid, we propose to decompose it into two structured components, which are then used as the basis for verification through $V(q_{i})$. Specifically:
Atomic Conditions, $\mathcal{P}(q_{i})=\{P_{1},P_{2},\dots,P_{N}\}$: These represent the foundational mathematical statements embedded within the question, serving as the building blocks for its logical structure.
Target Goals, $\mathcal{G}(q_{i})=\{G_{1},G_{2},\dots,G_{M}\}$: These define the specific objectives or queries that the question aims to address, encapsulating its intended purpose.
These two components are systematically combined in text form to reconstruct the complete question statement, enabling a rigorous evaluation of its correctness and logical coherence.
Figure 1: An overview of our MathQ-Verify framework. Given a math question, we extract atomic conditions $P_{i}$ and the goal $G_{i}$, and conduct a five-step verification process: (1) detection of contaminated instructions, (2) linguistic error detection, (3) atomic condition verification, (4) cross-condition conflict detection, and (5) condition completeness validation. Only questions that pass all checks are retained as correct structured math questions.
3.2 Question Correctness Verification
As outlined in the previous section, verifying the correctness of a mathematical question, denoted as $V(q_{i})$, requires a systematic approach to deconstruct and analyze its components. In this section, we introduce the five stages of the Math Question Verification (MathQ-Verify) pipeline, which is specifically designed for this purpose. Let’s now verify the math question step by step.
Step1: Contaminated Instruction Detection.
To construct a high-quality textual math QA dataset, the first step is to detect and remove contaminated instructions, such as those containing misleading instructions or leaked answers. This process ensures that only semantically and syntactically appropriate items are retained for downstream tasks. To this end, we define a binary instruction validity indicator for each question $q_{i}$ as:
$\text{InstValid}(q_{i})=1\iff\text{IsMathQ}(q_{i})\land\neg\text{HasMisleadingCue}(q_{i})\land\neg\text{HasAnswerLeak}(q_{i})$
This holds if and only if $q_{i}$ is a genuine math question, free from misleading linguistic patterns and explicit answer leakage. The indicator reflects whether the instruction adheres to the structural and semantic criteria expected of high-quality mathematical queries. In particular, we perform a model-centric inspection that assesses both the mathematical nature and instructional integrity of each input. For example, if a question $q_{i}$ includes textual cues such as “please rewrite” or “please rephrase”, or if it appends an explicit answer within the same prompt, it is considered misaligned with the problem-solving objective. Such artifacts may distract the model from focusing on solving the mathematical content and are thus filtered out during this validation phase.
Step2: Linguistic Error Detection.
After identifying and isolating the core mathematical question in the first step, we proceed to detect linguistic imperfections that may hinder readability or model interpretation. This step focuses on linguistic-level issues such as spelling mistakes, grammatical errors, and LaTeX formatting anomalies. While these can be addressed with rule-based heuristics, we adopt the Qwen-2.5-7B-Instruct model for simplicity and generality.
Following this, we obtain a set of clean math problems, where "clean" refers to questions that are exclusively mathematical in form, regardless of clarity or correctness. Formally, a question $q_{i}$ is clean if it contains no non-semantic errors that affect surface quality, and its content can be decomposed into a set of mathematical atomic conditions $\mathcal{P}(q_{i})$ and a target query $\mathrm{Query}(q_{i})$.
$\text{Clean}(q_{i})=1\iff\neg\,\text{LinguisticError}(q_{i})\land\exists\,\mathcal{P}(q_{i}),\,\mathrm{Query}(q_{i})$
Step3: Atomic Condition Error Detection.
For a clean mathematical question $q_{i}$, we systematically examine each of its atomic conditions $\{P_{1},P_{2},\ldots,P_{N}\}$ to ensure their consistency with established mathematical principles. A atomic condition is defined as a foundational statement that encapsulates a distinct mathematical fact (e.g., $x\in\mathbb{Z}$). Each condition $P_{j}$ undergoes rigorous validation to confirm its adherence to the rules of the corresponding mathematical domain. To guarantee validity, we verify the following for all conditions $P_{j}\in P(q_{i})$:
$\text{Valid}(q_{i})=1\iff\forall P_{j}\in\mathcal{P}(q_{i}),\quad\text{Valid}(P_{j})=1$
While some flexibility may be allowed for commonly understood concepts or heuristic steps (often referred to as "magic steps") that aid in mathematical reasoning, any condition that contradicts fundamental definitions is strictly rejected. For example, defining a discrete quantity as continuous would be considered invalid. To illustrate, consider the following question, which includes a atomic condition that violates core mathematical principles:
Figure 2: An example of an incorrect atomic condition: a negative area contradicts fundamental mathematical definitions.
Step4: Contradictions among Conditions.
To assess the internal consistency of a math question $q_{i}$, we examine all subsets of its atomic conditions $\mathcal{P}(q)$ (with subset size $|S|$, $2\leq|S|\leq N$), ensuring no logical contradictions arise when conditions are considered jointly. At this stage, each condition is individually valid, and pairwise consistency has been verified; this step ensures global coherence across the full set. Formally, we require:
[…]
Step5: Condition Completeness Evaluation.
After verifying the absence of contradictions among atomic conditions, we assume the condition space $\mathcal{P}(q_{i})$ is internally consistent. Under this assumption, if a target query $G_{i}\in\mathcal{G}(q_{i})$ remains unsolvable, it indicates that essential information is missing. We formalize this by checking whether each goal is logically derivable from the available conditions:
$\text{Complete}(q_{i})=1\iff\forall G_{i}\in\mathcal{G}(q_{i}),\quad\text{Derivable}(G_{i}\mid\mathcal{P}(q_{i}))=1$
Redundant but non-conflicting conditions are permitted, as they do not compromise derivability. This criterion ensures that the classification of a question as under-specified stems solely from genuine informational gaps, not from internal inconsistency.
Combining all verification steps, the full verification function is given by:
$\mathcal{V}(q_{i})=\text{InstValid}(q_{i})\land\text{Clean}(q_{i})\land\text{Valid}(q_{i})\land\text{ContradictionFree}(q_{i})\land\text{Complete}(q_{i})$
3.3 Multi-Model Voting Strategy
To enhance the robustness of condition verification, we propose a multi-model voting strategy that aggregates predictions from multiple independently trained models. Instead of relying on a single model’s output, this strategy leverages majority voting to reach a collective decision. Formally, we consider configurations denoted as $(n,k)$, where $n$ is the total number of models in the ensemble, and $k$ is the minimum number of models that must agree (i.e., output a positive decision) for the final prediction to be accepted.
Let $V_{j}(q_{i})\in\{0,1\}$ denote the binary verification output of the $j$-th model on query $q_{i}$. The final decision $V^{\text{final}}(q_{i})$ is given by:
$V^{\text{final}}(q_{i})=\begin{cases}1,&\text{if }\sum\limits_{j=1}^{n}V_{j}(q_{i})\geq k\\ 0,&\text{otherwise}\end{cases}$ (1)
This formulation provides a tunable mechanism to balance precision and recall by adjusting $k$ relative to $n$. By requiring agreement among a subset of models, the strategy mitigates individual model biases and reduces variance in predictions, thereby increasing the overall reliability of the verification process. The empirical effectiveness of this approach is demonstrated in Section 5.
Table 1: MathQ-Verify and baseline performance on MathClean-GSM8K and MathClean-MATH.
Model Method MathClean-GSM8K MathClean-MATH
Acc (%) F1 (%) Inv. (#) Acc (%) F1 (%) Inv. (#)
Qwen2.5-7B Baseline 67.75 74.02 9 71.33 76.13 20
MathQ-Verify 74.38 76.09 4 76.75 77.06 11
Qwen2.5-Math-7B Baseline 67.63 71.94 99 68.00 71.53 330
MathQ-Verify 59.50 68.02 183 56.08 61.39 543
Llama-3.1-8B Baseline 59.62 46.96 2 64.88 58.82 5
MathQ-Verify 65.62 72.55 22 65.50 72.42 4
DeepSeek-R1-Distill-Qwen-7B Baseline 62.38 59.16 37 65.25 63.61 214
MathQ-Verify 66.88 73.76 26 68.04 74.10 133
4 ValiMath: A Step-Tagged Extension of MathClean
While MathClean[19] provides a benchmark for validating math problem correctness, it primarily contains simpler, synthetic questions and lacks fine-grained annotations. To introduce a more diverse range of difficulties for comprehensive evaluation, we present ValiMath, a benchmark of 2,147 questions (1,299 correct, 848 incorrect) covering five diverse error types as shown in Table 5.
Our dataset encompasses a wide range of difficulty levels and diverse mathematical domain labels. The validity of the data is ensured through a rigorous filtering and validation process. We have reorganized and developed a new secondary classification system based on the MSC classification standard [3], which has been systematically applied to our data. The dataset is constructed through a meticulous multi-stage pipeline, involving mathematics domain experts to guarantee high-quality annotations. This comprehensive process ensures both the correctness of the questions and the accuracy of the labels. More details and the complete data construction process, along with data analysis results, are provided in Appendix A.
5 Experiments
In this section, we conduct a series of rigorous experiments to systematically evaluate the effectiveness and impact of our proposed pipeline. Specifically, we aim to address the following research questions:
Q1: How does MathQ-Verify perform on established benchmarks, and is the proposed pipeline adaptable across different model architectures?
Q2: To what extent does majority voting enhance precision, thereby increasing the proportion of correctly identified questions within the filtered dataset?
Q3: Does the filtering mechanism employed by our pipeline introduce substantial distributional shifts in the resulting dataset?
5.1 Experimental Setup
Models.
We evaluate a total of 14 large language models in this study, encompassing both reasoning-oriented and non-reasoning models. The reasoning-oriented models, such as DeepSeek-R1-671B, Gemini-2.5-Pro, QwQ-32B and GPT-o4-mini, are explicitly optimized for handling complex tasks that require multi-step inferential reasoning. In contrast, the non-reasoning models, including Qwen2.5-7B, Qwen2.5-Math-7B, Qwen2.5-72B, Qwen2.5-Math-72B, Llama-3.1-8B, S1.1-32B, Llama-3.1-70B, and GPT-4o, are primarily designed for general-purpose language understanding or specialized mathematical reasoning without explicit support for extended reasoning chains.
Dataset.
We utilize the synthesized and annotated versions of GSM8K and MATH from the MathClean benchmark as our primary evaluation datasets. Additionally, we incorporate the ValiMath dataset, as described in Section 4, to comprehensively assess the performance of our approach.
Table 2: Comparison of our MathQ-Verify and baseline performance on ValiMath.
Model Method Acc F1 Prec Inv. S1&S2 S3 S4 S5
Non-reasoning Models
Qwen2.5-7B Baseline 67.49 76.08 68.56 1 73.64 77.32 69.54 73.22
Ours 67.91 77.61 67.15 4 78.71 81.56 75.83 80.25
Qwen2.5-Math-7B Baseline 64.14 71.92 68.33 377 63.48 67.63 68.05 66.60
Ours 60.18 67.77 66.40 615 60.92 62.93 63.81 61.71
Qwen2.5-72B Baseline 70.28 78.65 69.57 1 73.92 80.39 75.41 76.57
Ours 72.38 80.77 69.79 0 81.60 83.98 75.55 80.53
Qwen2.5-Math-72B Baseline 68.28 76.70 69.03 75 68.37 77.36 74.90 74.57
Ours 68.37 77.39 68.19 44 73.40 80.25 76.29 76.53
Llama-3.1-8B Baseline 59.52 70.87 62.77 2 72.47 75.87 71.36 75.13
Ours 63.25 73.43 65.27 41 72.66 76.06 71.91 75.97
Llama-3.1-70B Baseline 66.84 75.26 68.59 2 71.22 74.99 68.51 72.75
Ours 67.26 75.72 68.67 4 72.66 76.06 71.91 75.97
S1.1-32B Baseline 67.77 72.76 74.46 62 54.63 63.25 64.04 59.25
Ours 75.31 78.78 82.07 8 63.67 61.11 59.94 58.13
GPT-4o Baseline 66.60 71.74 73.51 13 58.97 62.83 58.87 58.92
Ours 72.29 77.52 76.11 6 70.33 66.65 60.18 63.48
Reasoning Models
GPT-o4-mini Baseline 73.31 77.59 78.86 0 58.22 64.42 65.21 61.25
Ours 79.23 83.36 80.88 4 73.82 68.84 64.32 65.21
Gemini-2.5-Pro Baseline 71.31 75.48 78.15 1 53.98 62.41 64.79 59.62
Ours 75.97 78.62 85.11 0 59.66 57.57 59.85 54.68
QwQ-32B Baseline 70.33 75.51 75.42 17 55.05 66.37 68.47 62.37
Ours 77.64 81.62 81.19 4 67.68 66.14 64.23 63.06
DeepSeek-R1-671B Baseline 69.40 73.13 78.01 0 49.74 60.13 62.69 56.96
Ours 74.80 76.79 86.72 0 56.82 54.82 56.36 51.00
Qwen3-235B-A22B Baseline 69.21 75.40 72.93 0 62.18 69.77 65.16 66.23
Ours 78.16 81.53 83.40 0 66.09 63.06 62.27 60.18
Evaluation Metrics.
We adopt standard evaluation metrics to assess model performance, including accuracy, precision, recall, F1 score, the number of invalid outputs, and step-wise accuracy. Accuracy reflects the overall correctness, while precision and recall measure the proportion of correct predictions among all valid outputs and among all answerable questions, respectively. The F1 score provides a balanced summary of these two metrics. Invalid outputs refer to the number of samples from which no valid answer can be extracted using either instruction-based or rule-based decoding; final judgments on invalidity are made through semantic evaluation using GPT-4o. Step-wise accuracy measures the correctness of each individual step executed in the MathQ-Verify framework. S1&2 to S5 represent the accuracy of predictions for Question correctness at each steps in the ValiMath. Notably, under comparable recall, our method consistently achieves higher precision, indicating stronger robustness in identifying valid questions and resisting spurious outputs.
Baseline.
For the analysis of invalid samples on the MathClean-Benchmark, we adhere to the evaluation methodology established in the original work. For the ValiMath dataset, we construct a baseline method that assesses the correctness of each input question directly, without employing the decomposition or multi-step verification procedures integral to the MathQ-Verify framework. This baseline provides a point of comparison to evaluate the efficacy of our proposed pipeline.
Settings.
All open-source models are deployed using the vLLM framework, while closed-source models are accessed via their respective official APIs. For open-source models, we adopt consistent settings: temperature=0.7, top_p=0.8, top_k=20, and max_tokens=16384. For math-specialized models (e.g., Qwen2.5-Math-7B), we set max_tokens=4096 to optimize performance. All experiments are conducted on 8 x NVIDIA H200 GPUs.
5.2 Results and Analysis
To address Q1, which investigates the effectiveness of the proposed MathQ-Verify framework, we conduct comprehensive evaluations on three benchmark datasets: MathClean-GSM8K, MathClean-MATH, and ValiMath. The primary objective is to quantify the extent to which MathQ-Verify improves prediction accuracy and to assess its applicability across diverse model architectures. As presented in Table 1, our method yields substantial improvements in both accuracy and the reduction of invalid outputs across a wide range of models. In particular, models such as Qwen2.5-7B and LLaMA-3.1-8B exhibit notable gains in precision and F1 scores. However, certain specialized models such as Qwen2.5-Math-7B demonstrate limitations in instruction-following capabilities, resulting in increased invalid outputs. These observations suggest that models with sufficient instruction-following capabilities generally exhibit performance improvements when augmented with MathQ-Verify.
To further evaluate the robustness and generalizability of MathQ-Verify, we extend our analysis beyond the MathClean benchmarks to the ValiMath dataset. This investigation aims to determine whether the observed performance gains persist under distributional shifts and dataset variations. As illustrated in Table 2, MathQ-Verify consistently enhances performance across datasets. For example, GPT-4o and S1.1-32B achieve an improvement of 5.78 and 6.02 percentage points in F1 score, respectively, alongside increases in precision. These findings affirm the framework’s ability to generalize across diverse data distributions, effectively improving both the reliability and precision of model outputs while mitigating spurious generations.
Furthermore, we conducted a stepwise evaluation from S1 to S5 to assess the contribution of each step to overall performance. This evaluation aimed to understand how the various stages of the pipeline interact with different model and how they influence performance. The results indicate that both non-reasoning and reasoning models exhibit improvements in the early stages (S1&2 and S3), with notable gains observed in models such as Qwen2.5-7B. However, reasoning models, such as QwQ-32B, showed a decline in accuracy during the later stages (S4 and S5). This step-wise analysis provides insights into how MathQ-Verify interacts with pipeline stages and reveals that its effectiveness is influenced by the model’s reasoning capabilities and task complexity.
Figure 3: Distribution heatmaps of correct questions across difficulty levels and primary mathematical categories. The left heatmap shows human-annotated correct question distribution, while the right heatmap shows the distribution of questions validated as correct by MathQ-Verify after filtering.
5.3 Multi-Model Voting Ensemble Trade-Offs.
In this subsection, we aim to address Q2, whether majority voting can further enhance precision. The numbers correspond to the models as indicated in Table 2.
To prioritize precision, we observe that increasing the voting threshold ($k$) leads to more conservative predictions. For instance, in the $(3,3)$ configuration, where unanimous agreement is required, precision peaks at 91.42%. However, this comes with a substantial drop in recall (61.51%), indicating that many potentially valid samples are filtered out.
Conversely, to maintain recall, configurations with lower $k$ values, such as $(3,1)$, offer a more balanced trade-off. This configuration achieves the highest F1 score (82.48%) alongside a recall of 86.99%, demonstrating its effectiveness in preserving a greater portion of correct samples while still enhancing prediction quality, as shown in Table 3.
Considering the trade-offs between precision and recall, and the objective of retaining a broad and diverse set of high-quality samples, the $(2,2)$ configuration emerges as the optimal choice. It provides an effective balance, enhancing mathematical problem verification while ensuring that recall remains acceptable. Different configurations support varying precision and cost needs. This approach highlights the flexibility of $(n,k)$ voting, enabling us to prioritize precision without sacrificing recall, and ultimately improving the overall performance of the pipeline.
Table 3: Different model ensembles for $(n,k)$ configurations.
$(n,k)$ Voters Precision (%) Recall (%) F1 (%) TP FP
(1, 1) 86.72 68.90 76.79 895 137
(2, 1) ① ② 83.19 79.21 81.15 1029 208
(2, 2) ① ② 89.56 62.74 73.79 815 95
(3, 1) ① ② ③ 78.42 86.99 82.48 1130 311
(3, 2) ① ② ③ 86.20 75.98 80.77 987 158
(3, 3) ① ② ④ 91.42 61.51 73.54 799 75
(4, 4) ① ② ④ ⑤ 92.19 53.58 67.77 696 59
(5, 5) ① ② ④ ⑤ ⑥ 92.44 50.81 65.57 660 54
5.4 Consistency of Data Distribution
To answer Q3, which examines whether MathQ-Verify introduces significant distributional bias, we compare the joint category–difficulty distributions of the original human-annotated correct questions with those filtered by MathQ-Verify, as shown in Figure 3. The heatmaps reveal that the filtered distribution closely matches the original, with only minor shifts observed in certain categories and difficulty levels. Low-density areas, such as extreme difficulty levels and peripheral categories, remain consistent in both distributions. These results indicate that MathQ-Verify preserves the original category–difficulty structure, maintaining consistency with the human-annotated distribution.
5.5 Ablation Study
To evaluate each step in the MathQ-Verify pipeline, we perform an ablation study by removing one verification module at a time. Table 5 reports results with GPT-o4-mini, where precision and F1 show the effect of each removal. Omitting the first two validation steps causes the largest drop—precision and F1 decrease by over 6% and 3% respectively. Removing contradiction detection slightly lowers performance by about 2% in precision. Removing condition completeness increases F1 but slightly reduces precision, indicating a trade-off between strictness and coverage. Each module contributes uniquely, and their combination achieves the best precision–recall balance.
Table 4: Data distribution of ValiMath.
Category Count %
Instruction Error 143 6.66
Nonsemantic Check 119 5.54
Minimal Domain Check 171 7.96
Contradiction Check 300 13.97
Completeness Check 115 5.36
Correct Instances 1299 60.50
Table 5: Ablation study.
Variant Precision (%) F1 (%)
Baseline 78.86 77.59
MathQ-Verify 80.88 $\uparrow 2.02$ 83.36 $\uparrow 5.77$
w/o Step 1 78.41 $\downarrow 2.47$ 81.81 $\downarrow 1.55$
w/o Step 2 74.50 $\downarrow 6.38$ 80.00 $\downarrow 3.36$
w/o Step 3 80.19 $\downarrow 0.69$ 82.84 $\downarrow 0.52$
w/o Step 4 78.91 $\downarrow 1.97$ 83.13 $\downarrow 0.23$
w/o Step 5 79.53 $\downarrow 1.35$ 83.93 $\uparrow 0.57$
6 Conclusion
In this paper, we introduce a high-precision filtering pipeline for mathematical questions in LLM training data, in contrast to prior work that predominantly filters based on answer correctness. First, we perform format-level, linguistic filtering to isolate well-formed mathematical questions. Next, we formalize each problem by defining atomic conditions and inter-condition consistency checks such as domain conformity and contradiction detection allowing us to decompose each question into verifiable units. Finally, a goal-driven completeness assessment rigorously identifies any missing conditions. To facilitate fine-grained, step-wise validation, we also construct the MVQ-2K benchmark, comprising 2147 human-verified examples. Experiments on different LLMs show that a single MathQ-Verify model achieves state-of-the-art performance on MathClean and MVQ-2K, and that merely a majority-voting method raises precision up to 92%. Our pipeline thus provides a robust, high-accuracy solution for constructing quality mathematical QA datasets.
Appendix A The Construction Process of ValiMath
Figure 4: Overview of the ValiMath Construction Pipeline
Synthesis.
For synthetic data in the mathematics domain, especially for the generation of questions, there is a high demand for diversity. Additionally, the synthetic data itself tends to align more closely with the distribution of model training data. To construct a diverse set of questions, while also introducing more challenging problems compared to MathClean [19], we thoroughly researched existing prompts used for generating synthetic questions. Based on our findings, we carefully designed a prompt that ensures the diversity of the generated questions. To ensure the generated mathematical problems possess a certain level of difficulty and challenge, we ultimately used NuminaMath [18] as our source data for question synthesis.
Furthermore, to capture a broad range of potential errors that the model might generate, we employed models of different sizes for question synthesis. Stronger reasoning models tend to generate problems with a lower error rate during normal question synthesis, while smaller models are more likely to produce erroneous questions. Therefore, we used a variety of models, including: Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, GPT-4o, and GPT-o3mini, ensuring that each model generated a similar number of questions. In total, we synthesized 10,000 mathematical questions.
Filtering.
To ensure the uniqueness and determinacy of the final answers, we applied a GPT-4o-based prompt filtering process to exclude all multiple-choice questions. It is important to note that, since our objective is to construct a dataset containing both correct and incorrect problems, no additional filtering was performed beyond this step. This preserves the natural distribution of model-generated outputs.
Data Selection.
To reduce annotation costs in advance, we observed during preliminary experiments that general-purpose models with strong reasoning ability—such as GPT-4o tend to generate more correct than incorrect mathematical problems in our raw data (with fewer than 30% incorrect samples). To address this, we first performed a coarse-grained filtering step to retain a higher proportion of incorrect problems for annotation.
Specifically, we employed Qwen2.5-72B-Instruct to directly assess the correctness of each generated problem using the baseline judgment method described in the main paper. Based on these results, we curated a subset of approximately 4,000 mathematical problems with a target ratio of correct to incorrect samples set at 1:3. This step effectively reduced downstream annotation workload while preserving as many erroneous problems as possible, thereby enhancing the effectiveness of subsequent data verification.
Expert Annotation.
After obtaining the initial set of 4,000 mathematical problems, we invited domain experts in mathematics to perform detailed annotations. The goal was to ensure the accuracy of the binary correctness labels ("correct" or "incorrect") and, for incorrect problems, to further annotate the specific steps where errors occurred, following the five-stage framework defined in MathQ-Verify. The annotation process proceeded as follows:
  1. 1.
    Annotation Guidelines. Experts were provided with comprehensive annotation instructions, including precise definitions of each step in the MathQ-Verify process as outlined in the main paper. Each problem was evaluated across the five reasoning stages, and labeled accordingly as either “correct” or “incorrect.” For incorrect problems, annotators further specified the earliest step at which the error occurred.
  2. 2.
    Independent Dual Annotation. Each problem was independently annotated by at least two domain experts to mitigate individual biases. Discrepancies between annotations were resolved through a final joint review process, during which experts reached a consensus based on predefined labeling criteria.
  3. 3.
    Cross-validation. Following the dual annotation stage, we conducted a round of cross-validation to verify the reliability of the annotations. This involved a secondary review of previously labeled data, focusing on samples with disagreement, to ensure consistency and correctness in the final labels.
  4. 4.
    Subjectivity Analysis. To further assess and control for annotator subjectivity, we randomly sampled a subset of annotated problems for inter-annotator agreement analysis. We applied standard consistency metrics to quantify labeling stability across experts and identify any systemic annotation biases.
  5. 5.
    Quality Assurance. Through the above multi-step annotation workflow, we ensured the reliability, consistency, and accuracy of the labeled dataset, providing a high-quality foundation for downstream evaluation and model training.
Appendix B Additional Implement Details
B.1 Question Labels
We use GPT-4o to add the following question labels:
(1) Categorization: We adopt the Mathematics Subject Classification (MSC) taxonomy [3], which provides an extensive set of categories for mathematical topics. Given the complexity and large number of categories within MSC, we collaborated with domain experts to create a refined hierarchical classification scheme. The original categories were condensed into 7 primary categories, ensuring clear, non-overlapping coverage, and we introduced several secondary categories to capture more specific distinctions.
(2) Difficulty Classification: To classify the difficulty level of mathematical questions, we use the method of OMNI-MATH [13]. Since shot-level supervision was insufficient, we manually curated specific representative exemplars to ensure accurate delineation of difficulty levels for the dataset. This manual curation enabled us to better define the boundaries between different difficulty levels, facilitating more reliable analysis.
B.2 Handling the Invalid response from model
Although we explicitly specify the required output format during generation, models of varying sizes, capabilities, and contextual understanding do not always conform to the expected structure. In practice, most models produce only a small number of format violations, which can still be reliably handled via regular expressions to extract the final answers.
However, for math-specialized models, it is normal to observe a significantly higher incidence of unstructured or malformed outputs. These outputs often deviate from the expected format in ways that are not easily recoverable through rule-based methods, making it difficult to extract the final answer even with robust regular expression patterns.
To address this, we incorporate GPT-4o as a semantic parser to interpret and resolve ambiguous or non-compliant outputs. Specifically, for any model response that fails standard parsing, GPT-4o is prompted to infer the intended final answer based on the full output context. This ensures that downstream evaluation remains consistent and robust across diverse model types.